UI LABI 2 4oom 2.2504 405 m2 27) If 850 mL of 1.36 M NH.Cl is diluted to 4.50 L, the concentration of the resulting solution help with 27-29 please
Answer #1

27)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 1.36 M

V1 = 0.85 L

V2 = 4.5 L

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (1.36*0.85)/4.5

M2 = 0.257 M

Answer: E

28)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 0.75 M

M2 = 0.25 M

V2 = 25 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (0.25*25)/0.75

V1 = 8.3 mL

Answer: E

Only 1 question at a time please